OpenStudy (ajanijones):

medallss 60% of a number is 300. What is 25% of the number?

OpenStudy (anonymous):

0.60 * 300 = answer answer * 0.25 = second answer

OpenStudy (anonymous):

I don't think that's correct @DDS

OpenStudy (anonymous):

I just learned this ! @dss please explain the answer next time

OpenStudy (anonymous):

It's pretty explanatory

OpenStudy (anonymous):

of in math means to multiply

OpenStudy (ajanijones):

125 B. 45 C. 255 D. 720

OpenStudy (ajanijones):

thats my only anwsers

OpenStudy (anonymous):

0.60 * 300 = 180 180 * 0.25 = 45

OpenStudy (anonymous):

because if 60% of "a number" is equal to 300 then...\[x \times \frac{ 60 }{ 100 } = 300\]

OpenStudy (anonymous):

so X would be 500

OpenStudy (anonymous):

I would say 75 percent means hundredths. so 25 percent means 25 hundredths. so your question can also be written as "what is 25 hundredths of 300?". Well. One hundredth of 300 is 300 divided by 100....OK? which is 3........OK still? so 25 hundredths (or 25 percent) of 300 is 25 times 3 which equals 75.

OpenStudy (anonymous):

Bad wolf, you are right....300/.60

OpenStudy (anonymous):

and\[500 \times \frac{ 25 }{ 100 } =\]

OpenStudy (anonymous):

500 * 0.25 = 125

OpenStudy (anonymous):

@TheBadWolf you would be right i think but yah ....

OpenStudy (anonymous):

making the answer A

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I misread the problem originally

OpenStudy (anonymous):

How is it that all of your posts are getting medals instantly?

OpenStudy (anonymous):

I'm going with A or b

OpenStudy (anonymous):

oh nevermind i get it